# cp3-bench: A tool for benchmarking symbolic regression algorithms tested with cosmology

### v1

@@ 2.3 Cosmic backreaction and two-region models
where to reach the second line we simply use that the empty region is coasting and hence $q_{u}=0$. The last term is the backreaction term due to $Q_{D}$. By comparing with the Equations 2.6 we see that we can write $Q_{D}$ as e.g.
@@
The above two expressions are fairly straightforward analytical expressions. However, both expressions contain several dependent variables since $v,h,H_{D},\ddot{a}_{D}$ and $a_{D}$ all depend on time (differently) which may make the regression task more complicated. We will therefore include $\left(Q_{D},H_{D},v,h\right)$ and $\left(Q_{D},H_{D_{0}},\Omega_{m,0},a_{D},\ddot{a}_{D}\right)$ as datasets used for our benchmark to complement the earlier datasets for the target $Q_{D}$ where there is no known ground truth expression.
@@ 2.3.1 Redshift drift in 2-region models
The redshift drift is a particularly interesting observable in relation to backreaction and the backreaction conjecture because it can be used as a smoking gun signal for distinguishing between accelerated expansion due to dark energy versus apparent acceleration due to backreaction [smokinggun1, smokinggun2, smokinggun3, smokinggun4]. The foundation for the smoking gun signal is that the mean redshift drift is not in general equal to the drift of the mean redshift in an inhomogeneous spacetime even if that spacetime is statistically homogeneous (where “mean” observations here indicate observations after taking the mean over many lines of sight in order to remove statistical fluctuations). Mathematically, we can write this statement as
@@
The procedure for computing the redshift drift in two-region models is to arrange copies of the two FLRW regions consecutively after one another along a light ray propagated according to the equations
@@
where the scale factor is evaluated according to which local FLRW model it is in at any given position along the light ray. The redshift $z$ is the local redshift along a light ray which approximates $\expectationvalue{z}$ as discussed in [smokinggun1]. The resulting $z$ and $\delta z$ data only approximates their mean counterparts as they do contain small statistical fluctuations which means that these datasets have a small statistical error applied to them.
By solving the above equations, we generate datasets of the form $(\expectationvalue{\delta z}/\delta t_{0},\expectationvalue{z},\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0})$, $(\expectationvalue{\delta z}/\delta t_{0},\expectationvalue{z},H_{D}(\expectationvalue{z}),\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0})$ and $((\expectationvalue{\delta z}-\delta\expectationvalue{z})/\delta t_{0},\expectationvalue{z},\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0})$, where the two latter datasets are inspired by the fact that we know that the symbolic expression for the FLRW redshift drift requires either an equation of state parameter or Hubble parameter. The datasets were generated using 50 equidistant points of $f_{o}\in[0.1,0.25]$ and approximately 200 values of the redshift in the interval $z\in[0.1,1]$.
@@
As with the earlier two-region model datasets, we will assess the quality of the symbolic expressions obtained by the regression algorithms by seeing how well they generalized outside the model subspace the algorithms were trained on. Specifically, we make test datasets with the feature intervals $z\in[0.1,5]$ and $f_{o}\in[0.01,0.3]$.
@@ 2.4 Summary of data sets
Our cosmological datasets used for the benchmark are summarized in Table 1. Note that some targets are scaled to be of the order 1-10 in order to improve the symbolic regression. The feature intervals indicated in the table are for training data only. For datasets with unknown ground truth we assess the output from the symbolic regressions algorithms by comparing with a test dataset where the feature intervals are broadened. If the obtained symbolic expressions signify (the hitherto unknown) ground truths, they must perform well not only within the training interval but also in the test data representing the broadened parameter spaces. Otherwise, a symbolic expression with small MSE on the training data will simply represent a good (over-)fit without any (clear) physically relevant information. For the test data without known ground truths we use $z\in[0.1,5]$ and $f_{o}\in[0.01,0.3]$ to test the symbolic expressions outside the training regions.
Several datasets correspond to the same data but where different features are provided in the datasets. For some datasets, e.g. $C5$ and $C6$, this is because features used for parametrizing targets in cosmology are often inter-dependent. We wish to see how such inter-dependence affects the regression algorithms’ ability to find the ground truth expressions. Therefore we parameterize the same datasets using different parameters with some datasets having inter-dependent features and some datasets having strictly independent features. For other datasets, we use different parameterizations because the underlying ground truth is not know. We therefore do not know exactly which features the symbolic regression algorithms will benefit from knowing. For all the datasets based on two-region models we additionally note that the models are parameterized by a single parameter $f_{o}$ which is, however, not useful for obtaining general expressions for the targets valid also for cosmological settings not based on two-region models. Datasets based on two-region models are therefore considered both with $f_{o}$ as a feature and with more cosmologically generic features such as density parameters.
@@
Table 1: Summary of cosmological datasets used in the benchmark. In the rightmost column we indicate whether the given dataset has been attributed a Gaussian distributed 10 % error or no error (indicated by an ✗). The datasets based on redshift drift in two-region models have an un-quantified minor error which we indicate by (✗). $N$ indicates the number of equidistant points along each feature dimension. For the two-region redshift drift data, the data points are not exactly equidistant along the $z-$dimension. In addition, $N$ for the $z-$interval in these datasets are approximately 200 while the $N$ indicated in the table is for $f_{o}$ only. Datasets are indicated as $(\rm target|features)$.

### v2

@@ 2.3 Cosmic backreaction and two-region models
where the second line follows from the empty region being coasting (hence $q_{u}=0$). The last term is the backreaction term due to $Q_{D}$. By comparing with the Equations 2.6 we see that we can write $Q_{D}$ as e.g.
@@
The two above expressions for $Q_{D}$ are fairly simple analytical expressions. However, both expressions contain several dependent variables since $v,h,H_{D},\ddot{a}_{D}$ and $a_{D}$ all depend on time (differently) which may make the regression task more complicated. We will therefore include $\left(Q_{D},H_{D},v,h\right)$ and $\left(Q_{D},H_{D_{0}},\Omega_{m,0},a_{D},\ddot{a}_{D}\right)$ as datasets used for our benchmark to complement the earlier datasets for the target $Q_{D}$ where there is no known ground truth expression. The two datasets (denoted $C3g$ and $C3h$) are shown in Figure 4.
Figure 5: Top left: $\expectationvalue{\delta z}/\delta t_{0}$ plotted against $\expectationvalue{z}$, corresponding to $C4a$ and $C4b$. Top right: Scaled $\Delta z:=(\expectationvalue{\delta z}-\delta\expectationvalue{z})$ plotted against the features in dataset $C4e$. Bottom: $\expectationvalue{\delta z}/\delta t_{0}$ plotted against features in datasets $C4c$ and $C4d$.
@@ 2.3.1 Redshift drift in 2-region models
The redshift drift is a particularly interesting observable in relation to backreaction and the backreaction conjecture because it can be used as a smoking gun signal for distinguishing between accelerated expansion due to dark energy versus apparent acceleration due to backreaction [smokinggun1, smokinggun2, smokinggun3, smokinggun4]. The foundation for the smoking gun signal is that the mean redshift drift is not in general equal to the drift of the mean redshift in an inhomogeneous spacetime even if that spacetime is statistically homogeneous and isotropic (where “mean” observations here means observations after taking the mean over many lines of sight in order to remove statistical fluctuations). Mathematically, we can write this statement as
@@
The procedure for computing the redshift drift in two-region models is to arrange copies of the two FLRW regions consecutively after one another along a light ray. The regions must be arranged so that their proper distances appear in the same fraction as the proper volume fractions used for computing the volume averages (see e.g. section III a of [Bull_2012] for details). As discussed in [smokinggun1], each region must be small enough for them not to evolve much during the time it takes a light ray to traverse the homogeneity scale (set by the size of two consecutive regions). The regions were here chosen to have present-time size of order $10$Mpc. Small regions of proper length $\sim 10$Mpc will lead to small, rapid fluctuations around the mean. Larger regions would lead to large deviations from the mean which would imply that a single light ray could not capture the mean behavior of the redshift drift.
The light ray is propagated through the consecutive regions according to the equations
@@
where the scale factor is evaluated according to which local FLRW model the light ray is in at any given point along its path. The redshift $z$ is the local redshift along a light ray which approximates $\expectationvalue{z}$ as discussed in [smokinggun1]. The resulting $z$ and $\delta z$ data only approximate their mean counterparts since they contain statistical fluctuations as discussed above. The statistical fluctuations mean that these datasets have a statistical error applied to them. This can be seen in the left plot in Figure 5 which shows $C4a$ and $C4b$ plotted together. For a single value of $f_{o}$, the smoothed data that we wish to identify the ground truth for would correspond to a smooth/“best fit” line to the black squiggly data corresponding to $C4a$. It would be possible to provide smoothed data by computing the redshift and redshift drift along many (1001000$\sim$ 100-1000)∼ 100 - 1000 ) light rays and taking the mean. However, since real data contains similar fluctuations (as well as errors from imprecise measurements), we here choose to use the dataset reflecting the statistical fluctuations.
By solving the above equations, we generate datasets for a single two-region model with $f_{o}=0.2$. From this, we generate $C4a=(\expectationvalue{\delta z}/\delta t_{0},\expectationvalue{z})$. We then vary $f_{o}$ to obtain datasets of the form:
  • $C4b=(\expectationvalue{\delta z}/\delta t_{0},\expectationvalue{z},f_{o})$
  • $C4c=(\expectationvalue{\delta z}/\delta t_{0},\expectationvalue{z},\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0})$
  • $C4d=(\expectationvalue{\delta z}/\delta t_{0},\expectationvalue{z},H_{D}(\expectationvalue{z}),\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0})$
  • $C4e=((\expectationvalue{\delta z}-\delta\expectationvalue{z})/\delta t_{0},\expectationvalue{z},\Omega_{m,0},\Omega_{R,0},\Omega_{Q,0})$,
where the two latter datasets are inspired by the fact that we know that the symbolic expression for the FLRW redshift drift requires either an equation of state parameter or Hubble parameter. The datasets were generated using 50 equidistant points of $f_{o}\in[0.1,0.25]$ and approximately 200 values of the redshift in the interval $z\in[0,1]$ [^5] The exact number of redshift values in the dataset is difficult to control for these computations where we propagate light rays through an inhomogeneous universe and for this reason we chose, for convenience, to use equidistant points in the affine parameter rather than the redshift. This has the risk of leading to severe under-sampling in redshift space at high redshifts compared to lower redshifts. However, as long as we do not go further than $\sim z=1$ in the training sample, the distribution in redshift space stays fairly uniform.. Note in particular that $C4c$ and $C4e$ have the same features and that the difference is thus in the target where we in $C4e$ have subtracted the naive generalisation of the FLRW redshift drift. The datasets are plotted in Figure 5. It is not apparent from these plots that we gain much from subtracting $\delta\expectationvalue{z}$ from the target in $C4e$, although the plots for $C4e$ may be a bit less curved than those for $C4c$. Nonetheless, there is good physical reason to expect that a significant part of the redshift drift signal is given by the naive counterpart, and we therefore retain $C4e$ as part of our benchmark dataset.
@@
As with the earlier two-region model datasets, we will assess the quality of the symbolic expressions obtained by the regression algorithms by seeing how well they generalized outside the parameter subspace that the algorithms were trained on. Specifically, we make test datasets with the feature intervals $z\in[0,5]$ and $f_{o}\in[0.01,0.3]$.
Figure 6: Datasets $C5a-e$. Note that the dataset $C5f$ would be a combination of the plots in the middle and right figure.
2.4 Dark matter halo profiles
Next after dark energy, the most dominant component in the Universe is dark matter (see e.g. [Lisanti_2016] for an introduction). A pressing task in astroparticle physics is determining the distribution of dark matter in large objects such as galaxies. Simulations based on the $\Lambda$CDM model suggest a cuspy matter density profile with an increasing density towards the core of the galaxy. Although galaxies with a cusp profile indeed have been observed, other observations of objects such as dwarf galaxies show a core type density distribution with a constant density near the core (for more details see e.g. [de_Blok_2009]). The density profiles are sensitive to the specific qualities of the dark matter particles such as whether or not they interact through other forces than gravity. The study [Nguyen_2020] for instance indicates that self-interactions of dark matter particles can lead to more core-like profiles within the $\Lambda$CDM model. Such interactions could thus be necessary for reconciling observations with model predictions of the dark matter density profiles. Overall, the physical origin of dark matter is still one of the biggest mysteries in fundamental physics and it is even not clear that dark matter represents new particle types and not e.g. a need for a modification of our theory of gravity – or a combination, as discussed (and challenged) in e.g. [Eriksen_2021, frandsen_2018]. Learning more about the dark matter halo profiles of galaxies is one possible step towards a better understanding of dark matter phenomenology. The observational appearance of different dark matter halo density profiles (cusp versus core) for instance prompt the question of whether it is possible to identify another underlying more general density profile that can reconcile the cuspy and core profiles discussed today. To assess the utility of symbolic regression for identifying such unknown general density profiles, we include synthetic dark matter halo density profiles in our benchmark.
We consider two dark matter halo profiles. One represents a core profile while the other represents the cusp profile. The Navarro-Frenk-White (NFW) density profile (Equation 3 of [Navarro_1996]) representing a cusp profile is given as
$\rho_{\rm NFW}\propto\frac{1}{\frac{r}{R_{0}}(1+r/R_{0})^{2}}$ (2.17)
and the density profile for a core profile is given as [Burkert_1995]
$\rho_{\rm core}\propto\frac{R_{0}^{3}}{(r+R_{0})(r^{2}+R_{0}^{2})},$ (2.18)
where $R_{0}$ is a characteristic radius. We will set the proportionality constant to $1$ in the above relations and construct six datasets from the resulting halo profiles. The first three datasets will be with constant $R_{0}=1$ and use the NFW profile where we sample the density on a uniform grid of $N=1000$ points with $r\in[0.1,2]$. The three datasets will differ by the first having no error and the others having 1% and 10% Gaussian noise, respectively. We will additionally add a classification parameter $x$ as an extra feature which we set to $1$ for the NFW profile and $-1$ for the core profile. The datasets thus take the form $C5a=(\rho_{\rm NFW},r,x)$ and similarly for $C5b$ and $C5c$. For the remaining three datasets we will vary $R_{0}\in[0.5,1.5]$ and $r\in[0.1,2]$ on uniform grids with $N=100$ in each feature dimension. We include no errors in these datasets. We make one dataset for each of the two density profiles, yielding $C5d=(\rho_{\rm NFW},r,R_{0},x=1)$ and $C5e=(\rho_{\rm core},r,R_{0},x=-1)$. Since the classification feature has no impact on the target value in the five datasets, the SR algorithms should ideally be able to ignore this feature. However, we also consider a sixth dataset which is made by combining $C5d$ and $C5e$ to see if the SR algorithms can produce an analytical expression that faithfully represents both profiles. The resulting dataset is $C5f=(\rho_{\rm NFW/core},r,R_{0},x)$, where
$\rho_{\rm NFW/core}=\frac{1}{2}\left(\rho_{\rm NFW}+\rho_{\rm core}\right)+\frac{x}{2}\left(\rho_{\rm NFW}-\rho_{\rm core}\right).$ (2.19)
For this dataset, the value of $x$ is crucial for the algorithms to be able to distinguish between the two profiles. Such an expression extrapolates between two different dark matter density profiles, although using a fabricated classification feature. In reality, the desire is to use other features in real dark matter halo catalogs such as the SPARC [^6] http://astroweb.cwru.edu/SPARC/ catalog to identify a single universal expression for dark matter halo profiles. Since we are here interested in a demonstration benchmark we prefer to consider synthetic datasets with known ground truths and will therefore not consider real dark matter density profile data.
The datasets $C5a-C5f$ are shown in Figure 6. In the figure to the left we show the datasets $C5a,C5b$ and $C5c$. The figure also shows $\rho_{\rm core}$ for the same value of $R_{0}$ for comparison. The two other figures show the datasets $C5d$ and $C5e$, and then $C5f$ is a combination of these two.
Figure 7: Datasets $C6a$, $C6b$ and $C6c$. The subfigure to the left shows the dataset $C6a$ which contains $t$ as its single feature. The middle subfigure shows $C6b$ where the target ($h_{+}$) is plotted against all four features. The figure to the right shows $C6c$ plotted against its two parameters $t$ and $m_{1}$.
2.5 Gravitational wave-inspired oscillating dataset
Several cosmological datasets represent oscillating behavior, with prime examples being matter power spectra and the power spectrum of the CMB (cosmic microwave background) temperature fluctuations. Another well-known example of oscillating data in astrophysics is that from gravitational waves. We would therefore like our demonstration benchmark to include a dataset exhibiting oscillating behavior. However, for this demonstration we prefer datasets with known ground truths since this makes it easier to evaluate the performance of the SR algorithms. We will therefore not consider genuine CMB power spectra, matter power spectrum data or gravitational wave data as these do not represent known analytical expressions we can easily write up. Instead, we will introduce data based on an analytical expression that approximates a (very) simplified gravitational wave signal emitted from the in-spiraling of a two-body system. We construct the model by simplifying the inspiral model presented in [Buskirk_2019]. In particular, we consider the plus polarization mode (setting $c=G=1$)
$h_{+}=-\frac{2M\eta}{R}\left[\left(-\dot{r}^{2}+r^{2}\dot{\Phi}^{2}+\frac{M}{r}\right)\cos(2\Phi)+2r\dot{r}\dot{\Phi}\sin(2\Phi)\right],$ (2.20)
where $R$ is the distance from the binary to the detector, $M=m_{1}+m_{2}$ is the total mass of the binary, $\eta=m_{1}m_{2}/M^{2}$, $r$ is the orbital separation between the two bodies, $\Phi$ is the phase of the gravitational wave signal, and dots denote time derivatives. This equation is supplemented by the auxiliary equations
$M\frac{d\Phi}{dt}=v^{3}$ (2.21)
and
$\frac{dv}{dt}=-\frac{M}{r^{2}},$ (2.22)
where $v$ is the orbital velocity. In [Buskirk_2019], the equation for $dv/dt$ is provided to several post-Newtonian orders, leading to numerical solutions for the gravitational waves. In order to best assess if some SR algorithms perform significantly better/worse on oscillating datasets than monotonic datasets, we here wish to construct datasets with known analytical ground truths. To obtain this we will construct mock gravitational wave data by making the (non-physical) ansatz
$v(t)=\left(v_{0}-3t\right)^{1/3},$ (2.23)
where we ignore units as we will eventually scale the target to take on values close to 1 anyway. Using this ansatz, we find
$\begin{split}M\dot{\Phi}=-3t+v_{0}\implies M\Phi=-\frac{3}{2}t^{2}+v_{0}t+\Phi_{0}.\end{split}$ (2.24)
Furthermore, from Kepler’s third law we have (see [Buskirk_2019]) $r=M/v^{2}$ from which we obtain
$\begin{split}r&=M\left(-3t+v_{0}\right)^{-2/3}\\ \dot{r}&=2M\left(-3t+v_{0}\right)^{-5/3}.\end{split}$ (2.25)
Using the above with $\Phi_{0}=0$, we can write
$\begin{split}h_{+}&=-\frac{2M\eta}{R}\Big{[}\left(-4M^{2}\left(-3t+v_{0}\right)^{-10/3}+2\left(-3t+v_{0}\right)^{2/3}\right)\cos\left(-3t^{2}/M+2v_{0}t/M\right)\\ &+4M\left((-3t+v_{0})\right)^{-4/3}\sin\left(-3t^{2}/M+v_{0}t/M\right)\Big{]}.\end{split}$ (2.26)
We do not claim that the resulting expression for $h_{+}$ is particularly realistic [^7] In fact, our ansatz for $v(t)$ is inconsistent with $r=M/v^{2}$. However, the ansatz leads to oscillating behavior of the type we are interested in here and since we are merely interested in mock data, we will not concern ourselves with the non-consistency of the solution., but it nonetheless provides us with an oscillating analytical expression inspired by genuine cosmological/astrophysical data. We specifically make 3 datasets based on this data. The first dataset will be for a single binary system where we fix $m_{1}=0.5=m_{2}$, $R_{0}=5$ and $v_{0}=10$. These values were chosen somewhat adhoc, keeping in mind that we prefer obtaining target values of the order close to $1$. We then make a dataset by choosing a uniform grid for $t\in[0,3]$ and $N=1000$ (we neglect units since we do not consider this a realistic model and all parameter values are chosen out of convenience rather than physical meaningfulness). We will consider both $C6a=(h_{+},t)$ and $C6b=(h_{+},r,\dot{r},\Phi,\dot{\Phi})$. While the latter expression is analytically simpler, the former expression corresponds more closely to a real signal, since time is one of the observables we can directly measure. The last dataset, $C6c$, is similar to $C6a$ but we now vary the extra model parameter $m_{1}$ so that the dataset becomes $C6c=\left(h_{+},t,m_{1}\right)$. For generating this dataset we choose $t\in[0,3]$ and $m_{1}\in[0.1,1.5]$ and $N=100$ in both dimensions.
The datasets in group $C6$ are shown in figure 7.
@@ 2.6 Summary of datasets
The cosmological datasets used for our benchmark are summarized in Table 1. Note that some targets are scaled to be of the order 1-10 in order to improve the symbolic regression. The feature intervals used for the benchmark are those indicated in the table. For datasets with unknown ground truth we assess the output from the symbolic regressions algorithms by comparing with a test dataset where the feature intervals are broadened. If the obtained symbolic expressions signify (the hitherto unknown) ground truths, they must perform well not only within the training interval but also in the test data representing the broadened parameter spaces. Otherwise, a symbolic expression with small mean square error (MSE) on the training data can simply represent a good (over-)fit without any (clear) physically relevant information. For the test data without known ground truths we use $z\in[0,5]$ and $f_{o}\in[0.01,0.3]$ to test the symbolic expressions outside the training regions.
Several datasets correspond to the same setup and to the same data but with different features provided in the datasets. We therefore split our data into several groups, with each group corresponding to a specific physical setup. For instance, the group $C1$ contains the datasets $C1a,C1b,C1c$ and $C1d$. Below, we list the dataset groups and summarize what questions we wish to answer/qualities we wish to demonstrate by using each group.
  • Group $C1$ consisting of datasets $C1a,C1b,C1c,C1d$: Dataset $C1a$ represents what should be a fairly simple regression task, where the target variable depends on a single feature with the ground truth being a simple rational function. We would thus expect most SR algorithms to be able to identify this expression. In $C1b$ we have added noise and by comparing the SR algorithms’ performances on $C1a$ versus $C1b$ we wish to demonstrate how much the added error affects this performance. The final datasets, $C1c$ and $C1d$, are the same as $C1a$ but augmented to include more cosmological models and hence represents datasets with more features. In $C1c$ we vary one model parameter while we vary two model parameters in $C1d$. By comparing the performance of SR algorithms on $C1a$ versus $C1c$ and $C1d$ we wish to demonstrate how much the dimension of the feature/parameter space affects the performance of the SR algorithms. Note that the extra parameter in $C1d$ compared to $C1c$ is $H_{0}$ which simply gives an overall scaling of $H(z)$ which means that the overall complexity of the final expression for the two datasets are very similar. This means that we can be fairly certain that any difference in performance on dataset $C1c$ versus $C1d$ is specifically related to the dimension of the feature space and not complexity of the ground truth.
  • Group $C2$ consisting of $C2a,C2b$: These two represent the redshift drift for a family of FLRW models. The two datasets represent the same data but in $C2a$ the target is parameterized by two features which are inter-dependent while in $C2b$ the features are independent. With these datasets we wish to demonstrate whether or not the interdependence of features is important for the performance of SR algorithms.
  • Group $C3$ consisting of $C3a-C3h$: These datasets all represent backreaction for the two-region toy-models. For $C3a-C3f$, the ground truth of the target ($Q_{D}$ or $R_{D}$) is unknown. We include these datasets in our demonstration benchmark partially to showcase datasets where SR could potentially be utilized to reveal a new physical relation. We also include this dataset to highlight the difficulties one faces when seeking to use SR to reveal ground truth relations. The first two datasets in this group, $C3a$ and $C3b$, represent the simpler situation where only a single two-region model is considered and hence where the target depends only on a single parameter/feature. The two datasets differ by their target since we do not have prior knowledge regarding whether it it easier to obtain a symbolic expressions for one or the other. The datasets $C3c-C3f$ correspond to data where we also vary the model parameter $f_{o}$ so that the datasets $C3c$ and $C3d$ both have two features and where, again, we make two datasets to have both $Q_{D}$ and $R_{D}$ as targets in case it turns out to be easier for the SR algorithms to obtain the ground truth expression for one over the other. Datasets $C3e$ and $C3f$ represent the same datasets (i.e. same model parameters and same redshift) but where we supply other features than $f_{o}$. Specifically, we introduce features similar to those used for FLRW spacetimes. Based on the dynamical equations in Equation 2.6 we expect that relations between the target and provided features exist. However, we note that one of the features in these two datasets, namely $H_{D_{0}}$ takes the same value at each data point in these two datasets. It will therefore be interesting to see if this parameter is included in the expressions reported by the SR algorithms since optimally this should not be the case. Finally, datasets $C3g$ and $C3h$ are included because the targets are here chosen such that we actually know the ground truth expression of the target in terms of the provided features. However, the features are inter-dependent which we suspect may compromise the SR algorithms’ ability to identify the ground truth expressions.
  • Group $C4$ consisting of $C4a-C4e$: As $C3$, this group of datasets represent unknown ground truths, this time of a direct observable (the redshift drift). We chose to introduce two groups of datasets representing ground truths to give the SR algorithms more opportunities to identify unknown ground truths. Dataset $C4a$ represents data for a single two-region model and hence contains only a single feature. The dataset $C4b$ represents data for a family of two-region models, parameterized by $f_{o}$. Datasets $C4c$ and $C4d$ represent the same data parameterized by different features. Based on the analytical considerations of e.g. [smokinggun3], we expect that is may be possible to express the target through some (or all) of the supplied features. The last dataset in this group, $C4e$, represents the same dataset as $C4d$, but where the target has been rewritten in terms of the FLRW limit of the general symbolic expression/ground truth we are seeking. All datasets in this group have a statistical error. This error comes from the method used for generating data, where we consider a single light ray in an inhomogeneous spacetime rather than smoothing over $\sim 100-1000$ light rays to remove statistical fluctuations. Because of this statistical error, we expect that the SR algorithms are more likely to identify ground truths for the dataset $C3$ (which has no error) than for $C4$.
  • Group $C5$ consisting of $C5a-C5e$: This group represents data with known ground truths. As demonstrated in $C3$ and $C4$, when we do not know the ground truth, we also do not know what parameters/features to add to our datasets. It is therefore likely that we will add features that are not relevant for the target. With this dataset, we wish to see to what extent the extra features affect the SR algorithms’ performance by adding the superfluous classification feature $x$ to the datasets $C5a-C5d$. To highlight the importance of data precision we also consider versions of $C5a$ with 1% and 10% Gaussian noise. We lastly add $C5e$ which is simply the combination of $C5a$ and $C5d$. The corresponding analytical expression of the target extrapolates between the two density profiles. Such an extrapolation is of interest for dark matter halo research and the task thus resembles a possible genuine SR task within this field.
  • Group $C6$ consisting of $C6a-C6c$. This dataset represents a simplified gravitational wave dataset which is first of all considered in order to assess how well the considered SR algorithms perform on datasets that exhibit oscillation compared to the more monotonous behavior of the other datasets we consider. $C6a$ contains a single feature, $t$, and represents a fairly complicated analytical expression. $C6b$ represents the same data points, but rather than having $t$ as feature, we introduced four features that are all functions of $t$ but which can be combined into a simpler analytical expression. This dataset is thus similar to $C2a$ where the two features are interdependent but combine to form the target through a relatively simple analytical expression. Lastly, $C6c$ is similar to the dataset $C6a$ except that we now vary an extra model parameter ($m_{1}$) to see how much the extra feature affects the performance of the SR algorithms.
@@
Table 1: Summary of cosmological datasets used in the demonstration benchmark. In the rightmost column we indicate whether the given dataset has been attributed a Gaussian distributed 10 % error, 1% error or no error (indicated by an ✗). As discussed in Section 2.3.1, datasets based on redshift drift in two-region models have an un-quantified error which we indicate by (✗). $N$ indicates the number of equidistant points along each feature dimension (so for a $d$-dimensional dataset with $N$ points in each dimension, the total number of points in the dataset would be $N^{d}$). For the two-region redshift drift data, the data points are not exactly equidistant along the $z-$dimension. In addition, $N$ for the $z-$interval in these datasets are approximately 200 while the $N$ indicated in the table is for $f_{o}$ only. Datasets are indicated as $(\rm target|features)$. The range is always given for the parameter that was varied even though the selected features may be different. For instance, $C6b$ was obtained by varying $t$ but $t$ was not itself used as a feature in this dataset where instead we use the features $r,\dot{r},\Phi,\dot{\Phi}$.